OpenStudy (jacobbirddog):

PLEASE HELP MEH! Which equation could be used to solve this problem? The sum of two consecutive integers is 129. Find the integers. A. 2n = 129 B. 2n + 1 = 129 C. n + 1 = 129 D. 2n + 2 = 129

OpenStudy (anonymous):

ok 128/2 = 64 2 * n (64) + 1 = 129

OpenStudy (jacobbirddog):

how does that translate into one of these equations stated (im using a translator so if something weird comes up ignore it)

OpenStudy (anonymous):

ok I was looking at 2n + 1 = 129 so 129 - 1 = 128 128 / 2 = 64

OpenStudy (jacobbirddog):

okay thank you so much... btw im wondering if my translation sofware works ok please say Y if it does thanks :)

OpenStudy (anonymous):

Yes and your very welcome

OpenStudy (jacobbirddog):

Ty :) i fanned and medallioned you

OpenStudy (anonymous):

Same here

OpenStudy (jacobbirddog):

thanks!

OpenStudy (anonymous):

welcome :D

OpenStudy (jacobbirddog):

u like RQ? or fiab?

OpenStudy (anonymous):

What?

OpenStudy (jacobbirddog):

it's LAPFOX TRAX youtube thing

OpenStudy (jacobbirddog):

music mainly

OpenStudy (anonymous):

Idk what either of those are I like to stick with country

OpenStudy (jacobbirddog):

lol i used to live in texas

OpenStudy (anonymous):

im in florida but i go to this hillbilly church in the middle of nowhere

OpenStudy (jacobbirddog):

haha

OpenStudy (anonymous):

yeah lol

OpenStudy (jacobbirddog):

i live in cali :P

OpenStudy (anonymous):

omg i would die to live there

OpenStudy (jacobbirddog):

SF not south cal i wish south cali but lol nope

OpenStudy (anonymous):

Like california righ?

OpenStudy (jacobbirddog):

yep

OpenStudy (anonymous):

Ya awesome

OpenStudy (jacobbirddog):

electronica :P

OpenStudy (anonymous):

ok ill check it out. Do u go to church?

OpenStudy (jacobbirddog):

JW

OpenStudy (anonymous):

Is that the initials?

OpenStudy (jacobbirddog):

so yes i do...

OpenStudy (anonymous):

ok that's cool!

OpenStudy (jacobbirddog):

yes its the initials we have a website its called www.jw.org

OpenStudy (anonymous):

ok thats cool our church has one too, gbc.com grace baptist church

OpenStudy (jacobbirddog):

cool.

OpenStudy (jacobbirddog):

i have another problem lol i can't figure it out

OpenStudy (anonymous):

ok lay it on me

OpenStudy (jacobbirddog):

Which equation could be used to solve this problem? The sum of two consecutive odd integers is 112. Find the integers. A. 2n + 3 = 112 B. 2n + 2 = 112 C. 2n + 1 = 112 D. n + 2 = 112

OpenStudy (jacobbirddog):

i am translating this from japanese lol

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

2n + 2 = 112 i divided 112 by 2 and got 56 then i subtracted 56 minus 2 and got 54 and 54 is even so it will divide by 2

OpenStudy (jacobbirddog):

you're good at this

OpenStudy (anonymous):

Thanks

OpenStudy (jacobbirddog):

i have like 3 more out of my 100 i already did the other 95 but i can't figure out the last 3

OpenStudy (anonymous):

ok ill do these last three but I have to get back to my school afterwards

OpenStudy (jacobbirddog):

Solve the given number problem. The sum of three consecutive even integers is –72. What is the least of these numbers? A. –26 B. –22 C. –30 D. 24

OpenStudy (jacobbirddog):

Solve the given number problem. The sum of two consecutive integers is 51. What are the numbers? A. 24 and 25 B. 25 and 26 C. 26 and 27 D. 20 and 31

OpenStudy (jacobbirddog):

The sum of two consecutive even integers is –66. What are the numbers? A. –30 and –28 B. –32 and –30 C. –34 and –32 D. –30 and –36

OpenStudy (anonymous):

1. I think its -30 but not 100% sure. 2. 25 & 26 add up to 51 3. -32 & -34 add up to -66

OpenStudy (jacobbirddog):

tyyyyyy if you need english help or japanese... im ya guy

OpenStudy (anonymous):

Thanks I'll try to remember that. :D

OpenStudy (jacobbirddog):

ok!

OpenStudy (anonymous):

well i gtg... :(

OpenStudy (anonymous):

i gtg do math work of my own

OpenStudy (jacobbirddog):

i passed tyyyyy!!!! i got an 70!

OpenStudy (anonymous):

was i wrong?

OpenStudy (jacobbirddog):

if i missed one more question on my algebra i wouldve failed thank you so much!

OpenStudy (jacobbirddog):

*would've

OpenStudy (jacobbirddog):

failed.*

OpenStudy (jacobbirddog):

*If

OpenStudy (anonymous):

ya i wish i could've helped more

OpenStudy (jacobbirddog):

thank you so much have a nice day!

OpenStudy (anonymous):

Ya you too!

OpenStudy (jacobbirddog):

Bye!

OpenStudy (anonymous):

Bye!